# Breaking Language Barriers or Reinforcing Bias? A Study of Gender and Racial Disparities in Multilingual Contrastive Vision–Language Models

### v1

@@ Abstract
Multilingual vision–language models promise universal image–text retrieval, yet their social biases remain under-explored. We present the first systematic audit of three public multilingual CLIP checkpoints—M–CLIP, NLLB–CLIP, and CAPIVARA–CLIP—across ten languages that vary in resource availability and grammatical gender. Using balanced subsets of FairFace and the PATA stereotype suite in a zero-shot setting, we quantify race and gender bias and measure stereotype amplification. Contrary to the assumption that multilinguality mitigates bias, every model exhibits stronger gender bias than its English-only baseline. CAPIVARA–CLIP shows its largest biases precisely in the low-resource languages it targets, while the shared cross-lingual encoder of NLLB–CLIP transports English gender stereotypes into gender-neutral languages; loosely coupled encoders largely avoid this transfer. Highly gendered languages consistently magnify all measured bias types, but even gender-neutral languages remain vulnerable when cross-lingual weight sharing imports foreign stereotypes. Aggregated metrics conceal language-specific “hot spots,” underscoring the need for fine-grained, language-aware bias evaluation in future multilingual vision–language research.
[…]
Breaking Language Barriers or Reinforcing Bias? A Study of Gender and Racial Disparities in Multilingual Contrastive Vision–Language Models
@@ 1 Introduction
[…]
@@
[…]
To answer, we audit three public multilingual CLIP variants—M-CLIP, NLLB-CLIP, and CAPIVARA-CLIP—on ten languages spanning-resource levels (English, French, Spanish vs. Portuguese, Xhosa, Hindi), grammatical gender (Spanish, French vs. Turkish, Finnish), and writing systems. Using FairFace and the PATA stereotype suite in a zero-shot set-up, we quantify (i) max-skew for gender and race, and (ii) stereotype amplification for criminality, negativity, , and agency categories.
@@
The rest of the paper details the audit protocol (§3), reports findings (§4), discusses implications (§5), and concludes with limitations and future work (§76).
@@ 2.1 Bias in Vision–Language Models
OpenAI’s CLIP unveiled the promise of large-scale contrastive pre-training, but also surfaced entrenched social stereotypes inherited from web corpora. Early audits such as Hamidieh et al. (2024) introduce So-B-IT, a 374-term taxonomy that shows CLIP disproportionately associates Muslim, Black and immigrant identities with toxic prompts, tracing these patterns back to LAION-400M. Scaling alone does not guarantee fairness: Al Sahili et al. (2025) disentangle encoder width, dataset size and corpus composition, revealing that larger models can amplify gender and race bias when the data are imbalanced.
To mitigate such biases, researchers intervene at different points in the embedding pipeline. Seth et al. (2023) learn a lightweight additive residual on the image branch (DeAR), erasing protected-attribute information while preserving zero-shot accuracy. Conversely, Chuang et al. (2023) correct only the text branch using calibrated projections derived from biased prompts. A joint perspective is offered by Dehdashtian et al. (2024), who formulate debiasing in a reproducing-kernel Hilbert space, simultaneously aligning image and text representations and reducing training time by up to 10×10$\times$10 ×. Beyond English, Moreira et al. (2024) tackle a Portuguese CLIP variant (FairPIVARA), cutting four bias types by as much as $98\%$ without hurting accuracy, while Luo et al. (2024) curate a demographically annotated medical dataset and apply optimal-transport debiasing in a safety-critical domain.
@@ 2.2 Bias in Multilingual NLP and Large Language Models
The effect of multilingual training is mixed. Training one multilingual $2.6$B-parameter model instead of multiple monolingual models reduces bias on CrowS-Pairs and BBQ in Nie et al. (2024). Yet Levy et al. (2023) report that multilingual fine-tuning can amplify sentiment bias across race, religion and gender. On the mitigation side, Xu et al. (2025) cast debiasing as a multi-objective, multi-agent optimisation problem, lowering StereoSet and BBQ bias by up to $88\%$ with $<7\%$ utility loss, while Shirafuji et al. (2025) subtract a “bias vector” in parameter space, improving SEAT without hurting GLUE. In machine translation, Stanovsky et al. (2019)’s WinoMT benchmark remains a touchstone, documenting pervasive gender mistranslations across eight language pairs.
@@ 2.3 Cross-Modal and Cross-Lingual Perspectives
Although vision–language and multilingual NLP research address different modalities, they converge on several open challenges. First, benchmark parity remains elusive: CLIP studies are largely English-centric, while multilingual NLP lacks vision-grounded bias probes. Second, attribute transfer poses a tension between debiasing and alignment, since removing bias in one branch or language can impair cross-modal or cross-lingual performance, motivating joint approaches (e.g., Dehdashtian et al. 2024) and rigorous evaluation across languages (e.g., Levy et al. 2023). Third, low-resource coverage is critical, as the most severe biases tend to occur in the scarcest data regimes (see Moreira et al. 2024; Mitchell et al. 2025), underscoring the need for more culturally and linguistically diverse corpora.
@@ 3 Methodology
We ground our methodology in a clear, replicable audit framework, where each design decision is driven by both theoretical rigor and practical relevance. We begin by defining a unified embedding space that underpins all evaluated checkpoints, ensuring comparability across architectures and training regimes. Building on this foundation, we then systematically survey the components of our study—spanning model families, target languages, curated datasets, templated probing scenarios, and bias quantification metricsd—providing a transparent roadmap that both justifies our choices and facilitates reproducibility.
@@ 3.1 Embedding Preliminaries
Vision–language models project images and texts into a shared space in which semantic similarity can be computed with a dot product. Formally, a frozen vision transformer $f_{\!v}$ maps an image $x$ to an embedding $v=f_{\!v}(x)\in\mathbb{R}^{d}$; a language-specific text encoder $f_{\!t}$ turns a prompt $t$ into $u=f_{\!t}(t)$. We adopt the cosine similarity
@@
3.2 Model Checkpoints
We evaluate three publicly available, parameter-efficient adaptations of CLIP that enable cross-lingual image–text retrieval without retraining the full vision backbone.
M-CLIP
Carlsson et al. (2022). M-CLIP retains the ViT-B/32 vision tower from OpenAI CLIP and replaces the text encoder with a distilled XLM-R model covering 68 languages. Distillation is carried out via a cross-lingual teacher–student objective that minimizes mean-squared error between CLIP’s English text embeddings and their translations, using 3.4 M caption pairs (MS-COCO, GCC, VizWiz). This design requires updating only the text tower, yielding a compact multilingual extension that matches English CLIP on XTD-11 (R@10 $\approx$ 90 %) and outperforms prior multilingual baselines across all measured languages.
NLLB-CLIP
Visheratin (2023). NLLB-CLIP swaps CLIP’s original text encoder for Meta’s 3.3 B-parameter NLLB-200, while freezing the ViT-B/32 vision backbone. Fine-tuning uses 106 k LAION–COCO image–caption pairs whose English captions have been auto-translated into 201 languages. On Crossmodal-3600 (36 languages) it achieves R@1$\approx$ 43 %, and on XTD-200 it excels in very low-resource settings (e.g. Quechua, Māori). This recipe maximizes language coverage with minimal changes to the vision encoder.
CAPIVARA-CLIP
dos Santos et al. (2023). CAPIVARA-CLIP augments CC3M by (i) regenerating captions with BLIP-2, (ii) translating them into target languages (e.g. Portuguese, Hindi, Xhosa), and (iii) fine-tuning OpenCLIP (XLM-R B/32) using LoRA adapters under the LiT objective. The synthetic dataset comprises 3.3 M new captions in addition to the originals. Evaluated on Flickr30k and MS-COCO in pt/hi/xh, it yields +6–21 pp mean-recall improvements over the OpenCLIP baseline, making it well-suited for rapid, low-resource adaptation via lightweight adapter layers.
English baselines.
We retain OpenAI CLIP (ViT-L/14 and ViT-B/32) and OpenCLIP (ViT-B/32) in their original form to establish a ceiling for accuracy and a floor for fairness.
3.3 Language Partitions
low-resource :$:\{\text{pt},\text{xh},\text{hi}\},$ high-resource :$:\{\text{en},\text{es},\text{fr}\},$
gender-less :$:\{\text{tr},\text{fa},\text{fi}\},$ gender-rich :$:\{\text{sl},\text{es},\text{fr}\}.$
3.4 Datasets
FairFace.
[…]
Protected-Attribute Tag Association (PATA).
[…]
@@ 3.5 Bias-Probing Tasks
Our probes are template based so that no ground-truth labels are required; the model’s own ranking reveals its implicit associations.
@@ 3.6 Bias Metrics
$s_{G}(g,c)\;=\;\frac{1}{|D_{g}|}\sum_{I\in D_{g}}\cos\bigl{(}v(I),\,e(c)\bigr{)},$
@@
$\mathrm{Skew}(A,B)\;=\;\max\!\Bigl{(}\bigl{|}\tfrac{p_{A}-p_{B}}{p_{B}}\bigr{|},\;\;\bigl{|}\tfrac{p_{B}-p_{A}}{p_{A}}\bigr{|}\Bigr{)}.$
To obtain a single summary statistic, we average $\mathrm{Skew}(A,B)$ over all unordered pairs $\{A,B\}\subset G$. This Max Skew upper-bounds any one pair’s relative bias across the attribute $G$.
Corpus-level harm rates. Independently, we record the fraction of portraits whose top-1 predicted label falls into one of the “harm” categories:
[…]
This captures the overall prevalence of harmful misclassifications in the model’s outputs.
@@ 4 Results and Analysis
We report max–skew scores for gender and race; lower is fairer(Table 2).
@@
DS Mdl Sz $G_{\max}$ $R_{\max}$
cr co ag cr co ag
@@
NLLB B32 .07 .07 2.58 .17 .05 4.49
Capivara B32 1.00 .00 .44 .31 .02 2.74
PATA CLIP B/32 1.19 0.04 0.15 3.59 0.19 0.14
CLIP L/14 0.23 0.15 0.20 1.54 0.39 0.10
OpenCLIP B/32 1.86 0.31 0.07 0.69 0.19 0.11
mCLIP L/14 1.21 0.72 0.02 0.22 0.14 4.49
NLLB B/32 6.35 0.12 0.2 0.49 0.20 2.49
CAPIVARA B/32 0.24 0.92 0.17 0.27 0.23 1.38
Table 1: Maximum gender skew ($G_{\max}$) and mean race skew ($R_{\max}$) for crime (cr), communion (co), and agency (ag) biases across two evaluation sets (FairFace, PATA). Highest-skew entries per column and dataset are shown in bold.
Table 1 compares each multilingual checkpoint with its English-only counterpart on the same English test images. The baseline CLIP L/14 exhibits a gender–crime skew of $0.23$; replacing its text tower with the distilled XLM-R in mCLIP almost quadruples that value to $0.88$, and adding LoRA adapters in CAPIVARA lifts it to $1.00$. The encoder-swap strategy of NLLB-CLIP is still more problematic, reaching $2.58$, eleven times the baseline. Race biases follow a similar pattern on the communion axis: CLIP L/14 starts at $0.25$, mCLIP rises to $2.24$, CAPIVARA to $2.74$, and NLLB peaks at $4.49$. Agency stereotypes are more architecture-sensitive: mCLIP delivers the highest gender–agency skew ($0.73$), whereas NLLB keeps that particular score close to the baseline ($0.16$); the vision-backbone size plays only a minor role once the text encoder is multilingual.
@@
mclip Fairface $\max s_{G}^{c}$ 0.87 0.12 0.12 0.13 0.14 0.58 0.13 0.08 1.27 0.07
$\max s_{G}^{com}$ 0.31 0.13 0.24 0.12 0.29 0.14 0.14 0.17 0.45 0.33
$\max s_{G}^{ag}$ 0.00 0.08 0.08 0.07 0.27 0.94 0.09 0.12 0.05 0.37
$\max s_{R}^{c}$ 0.40 0.40 0.27 0.10 0.39 0.08 0.26 0.41 0.23 0.04

### v3

# Breaking Language Barriers or Reinforcing Bias? A Study of Gender and Racial Disparities in Multilingual CLIP

@@ Abstract
Multilingual vision–language models (VLMs) promise universal image–text retrieval, yet their social biases remain under‑explored. We perform the first systematic audit of four public multilingual CLIP variants—M‑CLIP, NLLB‑CLIP, CAPIVARA‑CLIP, and the debiased SigLIP‑2—covering ten languages that differ in resource availability and morphological gender marking. Using balanced subsets of FairFace and the PATA stereotype suite in a zero‑shot setting, we quantify race and gender bias and measure stereotype amplification. Contrary to the intuition that multilinguality mitigates bias, every model exhibits stronger gender skew than its English‑only baseline. CAPIVARA‑CLIP shows its largest biases precisely in the low‑resource languages it targets, while the shared encoder of NLLB‑CLIP and SigLIP‑2 transfers English gender stereotypes into gender‑neutral languages; loosely coupled encoders largely avoid this leakage. Although SigLIP‑2 reduces agency and communion skews, it inherits—and in caption‑sparse contexts (e.g., Xhosa) amplifies—the English anchor’s crime associations. Highly gendered languages consistently magnify all bias types, yet gender‑neutral languages remain vulnerable whenever cross‑lingual weight sharing imports foreign stereotypes. Aggregated metrics thus mask language‑specific “hot spots,” underscoring the need for fine‑grained, language‑aware bias evaluation in future multilingual VLM research.
Breaking Language Barriers or Reinforcing Bias? A Study of Gender and Racial Disparities in Multilingual CLIP
@@ 1 Introduction
Contrastive vision–language pre-training has propelled breakthroughs in image retrieval, captioning, and zero-shot recognition. OpenAI’s CLIP model, trained on $400$ M English image–text pairs, aligns textual and visual embeddings so well that a frozen encoder can now be dropped into commercial search engines, content moderation pipelines, and accessibility tools (Radford et al., 2021). Recent work extends this recipe to multilingual settings by (i) distilling CLIP’s text tower into smaller encoders (ii) replacing it with pretrained multilingual language models or (iii) re‑training from scratch on Web‑scale data, enabling cross‑modal search for 100+ languages on a single GPU(Carlsson et al., 2022; Visheratin, 2023a; dos Santos et al., 2023; Tschannen et al., 2025).
@@
    Does multilingual training dilute or amplify social bias when compared to English-only CLIP?
To answer, we audit four public multilingual CLIP variants—M-CLIP, NLLB-CLIP, CAPIVARA-CLIP, and SIGLIP2—on ten languages spanning-resource levels (English, French, Spanish vs. Portuguese, Xhosa, Hindi), grammatical gender (Spanish, French vs. Turkish, Finnish), and writing systems. Using FairFace and the PATA stereotype suite in a zero-shot set-up, we quantify (i) max-skew for gender and race, (ii) symmetric KL-divergence, and (iii) stereotype amplification for criminality, negativity, , and agency categories.
@@
  • We demonstrate that prevailing debiasing techniques fall short—especially in low‑caption languages—leaving crime‑related stereotypes largely intact and highlighting the need for more robust, language‑aware mitigation strategies.
The rest of the paper details the audit protocol (§3), reports findings (§4), discusses implications (§5), and concludes with limitations and future work (§Limitations6).
@@ 2.1 Bias in Vision–Language Models
OpenAI’s CLIP unveiled the promise of large‐scale contrastive pre-training, but also surfaced entrenched social stereotypes inherited from web corpora. Early audits such as Hamidieh et al. (2024) introduce So-B-IT, a 374-term taxonomy that shows CLIP disproportionately associates Muslim, Black and immigrant identities with toxic prompts, tracing these patterns back to LAION-400M. Scaling alone does not guarantee fairness: Al Sahili et al. (2025) disentangle encoder width, dataset size and corpus composition, revealing that larger models can amplify gender and race bias when the data are imbalanced.
To mitigate such biases, researchers intervene at different points in the embedding pipeline. Seth et al. (2023) learn a lightweight additive residual on the image branch (DeAR), erasing protected-attribute information while preserving zero-shot accuracy. Conversely, Chuang et al. (2023) correct only the text branch using calibrated projections derived from biased prompts. A joint perspective is offered by Dehdashtian et al. (2024), who formulate debiasing in a reproducing-kernel Hilbert space, simultaneously aligning image and text representations and reducing training time by up to $10\times$. Beyond English, Moreira et al. (2024) tackle a Portuguese CLIP variant (FairPIVARA), cutting four bias types by as much as $98\%$ without hurting accuracy, while Luo et al. (2024) curate a demographically annotated medical dataset and apply optimal-transport debiasing in a safety-critical domain. At Web scale, Tschannen et al. (2025) introduce SigLIP 2, which combines language filtering, active‐sample selection (ACID) and a multi-objective loss to reduce female representation bias from 35 % to 7 % and shrink agency skews on ImageNet without harming accuracy. Complementing these CLIP-centric studies, friedrich2025magbig examine multilingual text-to-image (T2I) generators and introduce MAGBIG, a controlled benchmark of 3,630 prompts across nine languages used to evaluate five multilingual T2I models; they find strong language-specific gender skews and show that ostensibly neutral formulations (e.g., indirect descriptions or German gender-star forms) often fail to remove bias and can degrade text–image alignment.
@@ 2.2 Bias in Multilingual NLP and Large Language Models
The effect of multilingual training is mixed. Training one multilingual 2.62.6B-parameter model instead of multiple monolingual models reduces bias on CrowS-Pairs and BBQ in Nie et al. (2024). Yet Levy et al. (2023) report that multilingual fine-tuning can amplify sentiment bias across race, religion and gender. On the mitigation side, Xu et al. (2025) cast debiasing as a multi-objective, multi-agent optimisation problem, lowering StereoSet and BBQ bias by up to $88\%$ with $7\%$ utility loss, while Shirafuji et al. (2025) subtract a “bias vector” in parameter space, improving SEAT without hurting GLUE. In machine translation, Stanovsky et al. (2019)’s WinoMT benchmark remains a touchstone, documenting pervasive gender mistranslations across eight language pairs.
@@ 2.3 Cross-Modal and Cross-Lingual Perspectives
Although vision–language and multilingual NLP research address different modalities, they converge on several open challenges. First, benchmark parity remains elusive: CLIP studies are largely English-centric, while multilingual NLP lacks vision-grounded bias probes. Notably, friedrich2025magbig fill part of this gap for generative vision by providing a multilingual, controlled T2I benchmark and demonstrating that switching the language of otherwise identical prompts can magnify gender stereotypes. Second, attribute transfer poses a tension between debiasing and alignment, since removing bias in one branch or language can impair cross-modal or cross-lingual performance, motivating joint approaches (e.g., Dehdashtian et al. 2024) and rigorous evaluation across languages (e.g., Levy et al. 2023). Third, low-resource coverage is critical, as the most severe biases tend to occur in the scarcest data regimes (see Moreira et al. 2024; Mitchell et al. 2025), underscoring the need for more culturally and linguistically diverse corpora.
@@ 3 Methodology
We ground our methodology in a clear, replicable audit framework, where each design decision is driven by both theoretical rigor and practical relevance. We begin by defining a unified embedding space that underpins all evaluated checkpoints, ensuring comparability across architectures and training regimes. Building on this foundation, we then systematically survey the components of our study—spanning model families, target languages, curated datasets, templated probing scenarios, and bias quantification metrics—providing a transparent roadmap that both justifies our choices and facilitates reproducibility. Full implementation details, are in Appendix A.
@@ 3.1 Embedding Preliminaries
Vision–language models project images and texts into a shared space in which semantic similarity can be computed with a dot product. Formally, a frozen vision transformer $f_{\!v}$ maps an image $x$ to an embedding $v=f_{\!v}(x)\in\mathbb{R}^{d}$; a language‐specific text encoder $f_{\!t}$ turns a prompt $t$ into $u=f_{\!t}(t)$. We adopt the cosine similarity
@@
3.2 Models and Training Recipes
All checkpoints keep the vision tower frozen and modify only the text side of CLIP:
  • M‑CLIP (Carlsson et al., 2022): XLM‑R base distilled from English CLIP on translated MS‑COCO and GCC captions (68 langs).
  • NLLB‑CLIP (Visheratin, 2023b): swaps in the 3.3 B NLLB‑200 encoder; fine‑tunes on 106 k LAION–COCO images translated to 201 languages.
  • CAPIVARA‑CLIP (dos Santos et al., 2023): LoRA adapters on OpenCLIP ViT‑B/32, trained with BLIP‑2 paraphrases and MT into {pt, hi, xh}.
  • SigLIP 2 (Tschannen et al., 2025): trained from scratch on 40 B WebLI pairs (10 masked‑token losses plus the ClipTheBias filter.
OpenAI CLIP (ViT‑L/14, ViT‑B/32) and OpenCLIP (ViT‑B/32) serve as English baselines. More details are in A.1.
3.3 Languages and Data
We probe ten languages grouped by resource level and morphology: high‑resource {en, es, fr}, low‑resource {pt, hi, xh}, gender‑neutral {tr, fa, fi}, gender‑rich {sl, es, fr}.
Images come from two balanced face corpora: (i) FairFace (7 race × 2 gender; 10 954 test images) and (ii) PATA (5 race × 2 gender; 4 934 test images).
@@ 3.4 Bias-Probing Tasks
Our probes are template based so that no ground-truth labels are required; the model’s own ranking reveals its implicit associations. Templates were machine‑translated with GPT‑o3 and human‑checked in E.
@@ 3.5 Bias Metrics
$sc_{G}(g,c)\;=\;\frac{1}{|D_{g}|}\sum_{I\in D_{g}}\cos\bigl(v(I),\,e(c)\bigr),$
@@
$\mathrm{MaxSkew}(A,B)\;=\;\max\!\Bigl(\bigl|\tfrac{p_{A}-p_{B}}{p_{B}}\bigr|,\;\;\bigl|\tfrac{p_{B}-p_{A}}{p_{A}}\bigr|\Bigr).$
To obtain a single summary statistic, we average $\mathrm{MaxSkew}(A,B)$ over all unordered pairs $\{A,B\}\subset G$. This Max Skew upper-bounds any one pair’s relative bias across the attribute $G$.
KL Divergence for Gender.
Gender is binary in both datasets, so we can treat the negative‑trait rate as a Bernoulli parameter. Let $p_{\!f}$ (resp. $p_{\!m}$) be the fraction of negative attributions for female (resp. male). Each gender defines $P_{f}=[1-p_{\!f},\,p_{\!f}]$ and $P_{m}=[1-p_{\!m},\,p_{\!m}]$. We report
$\mathrm{KL}(f\|m)$ $=(1-p_{\!f})\ln\frac{1-p_{\!f}}{1-p_{\!m}}+p_{\!f}\ln\frac{p_{\!f}}{p_{\!m}},$ (1a)
$\mathrm{SKL}$ $=\tfrac{1}{2}\bigl(\mathrm{KL}(f\|m)+\mathrm{KL}(m\|f)\bigr).$ (1b)
Corpus‑level Harm Rate.
Finally, we record the share of images whose top‑1 prediction is Criminal, Animal, or the negative pole of any social trait. Unlike the relative metrics above, this is an absolute error rate indicating how often the model emits overtly harmful content. Formal definitions and illustrative examples are in subsection A.4. All code, prompts and evaluation scripts will be released upon acceptance.
@@ 4 Results and Analysis
We report max–skew scores for gender and race (Tables 2 and 4) and KL divergence for gender (Table 3); lower values indicate greater fairness. We also report the corpus-level harm rate (Table 6).
@@
Table 1 compares each multilingual checkpoint with its English-only counterpart on the same English test images. The baseline CLIP L/14 exhibits a gender–crime skew of $0.23$; replacing its text tower with the distilled XLM-R in mCLIP nearly quadruples that value to $0.87$, and adding LoRA adapters in CAPIVARA lifts it to $1.00$. The encoder-swap strategy of NLLB-CLIP is still more problematic, reaching $2.58$, over ten times the baseline. SigLIP2, while lower overall, still registers a nontrivial gender–crime skew ($0.44$), suggesting that multilingual debiasing can reduce—but not eliminate—stereotypes.
Race biases follow a similar pattern on the communion axis: CLIP L/14 starts at $0.25$; mCLIP rises to $2.24$, CAPIVARA to $2.74$, and NLLB peaks at $4.49$. SigLIP2 performs more moderately ($0.11$) but does not lead on any dimension.
Agency stereotypes vary more sharply by architecture: mCLIP delivers the highest gender–agency skew ($0.40$), NLLB remains close to the baseline ($0.20$), and SigLIP2 nearly erases it (0.020.02)—a pattern consistent across evaluation sets.
SigLIP2 appears the least biased model overall, but the skew profiles are non-uniform: crime associations remain most persistent, and KL divergence analysis reveals hidden disparities behind aggregate scores.
DS Mdl Sz $\max s_{G}$ $\max s_{R}$
c com ag c com ag
@@
NLLB B/32 .07 .07 2.58 .17 .05 4.49
CAPIVARA B/32 1.00 .00 .44 .31 .02 2.74
SigLIP2 B/16 .44 .37 .02 .49 .11 .10
PATA CLIP B/32 1.19 .04 .15 3.59 .19 .14
CLIP L14 .23 .15 .20 1.54 .39 .10
OpenCLIP B/32 1.86 .31 .07 .69 .19 .11
mCLIP L/14 1.21 .72 .02 .22 .14 4.49
NLLB B/32 6.35 .12 .20 .49 .20 2.49
CAPIVARA B/32 .24 .92 .17 .27 .23 1.38
SigLIP2 B/16 1.66 .11 .00 2.65 .29 .09
Table 1: English bias: unilingual vs. multilingual models. Maximum gender skew ($\max s_{G}$) and mean race skew ($\max s_{R}$) for crime (c), communion (com) and agency (ag) on English test images from FairFace and PATA. Each multilingual checkpoint is compared against its English‑only counterpart with the same vision backbone; highest skew per column and dataset is highlighted in green for agency(ag), orange for communion(com), and red for crime(c) .
@@
mclip $F$ $\max s_{G}^{c}$ 0.87 0.12 0.12 0.13 0.14 0.58 0.13 0.08 1.27 0.07
$F$ $\max s_{G}^{com}$ 0.31 0.13 0.24 0.12 0.29 0.14 0.14 0.17 0.45 0.33
$F$ $\max s_{G}^{ag}$ 0.00 0.08 0.08 0.07 0.27 0.94 0.09 0.12 0.05 0.37
PT $\max s_{G}^{c}$ 0.72 0.12 0.02 0.15 0.23 1.29 0.23 0.13 0.22 0.08
PT $\max s_{G}^{com}$ 0.22 0.28 0.53 0.41 0.44 0.25 0.37 0.25 0.28 0.30
PT $\max s_{G}^{ag}$ 0.02 0.09 0.66 0.10 0.09 0.37 0.10 0.30 0.18 0.33
NLLB $F$ $\max s_{G}^{c}$ 0.07 0.30 0.59 0.26 0.06 0.13 0.17 2.12 0.11 0.02
$F$ $\max s_{G}^{com}$ 0.17 0.28 0.12 0.04 0.07 0.12 0.08 0.42 0.21 0.07
$F$ $\max s_{G}^{ag}$ 0.07 0.15 0.23 0.13 0.07 0.04 0.09 0.24 0.04 0.03
PT $\max s_{G}^{c}$ 0.12 0.10 0.03 0.48 0.28 0.01 0.02 0.39 0.32 0.19
PT $\max s_{G}^{com}$ 0.49 0.57 0.23 0.13 0.15 0.13 0.40 0.19 0.22 0.07
PT $\max s_{G}^{ag}$ 0.20 0.28 0.18 0.17 0.35 0.34 0.25 0.28 0.22 0.08
CAPIVARA $F$ $\max s_{G}^{c}$ 1.00 1.05 0.16 0.09 0.17 0.16 0.44 0.31 1.08 1.22
$F$ $\max s_{G}^{com}$ 0.31 0.53 0.63 0.33 0.19 0.12 0.32 0.13 0.25 1.13
$F$ $\max s_{G}^{ag}$ 0.00 0.01 0.66 0.04 1.06 0.25 0.25 0.11 0.37 0.37
PT $\max s_{G}^{c}$ 0.92 0.65 0.05 0.25 0.02 0.03 0.24 0.04 0.19 1.20
PT $\max s_{G}^{com}$ 0.27 0.16 0.20 0.20 0.24 0.14 0.18 0.04 0.35 0.15
PT $\max s_{G}^{ag}$ 0.17 0.15 0.05 0.01 0.52 0.53 0.23 0.43 0.14 0.32
SIGLIP2 $F$ $\max s_{G}^{c}$ 0.44 0.93 3.16 0.21 0.55 0.94 0.42 0.28 3.77 3.24
$F$ $\max s_{G}^{com}$ 0.37 0.33 0.02 0.04 0.01 0.13 0.31 0.16 0.17 0.03
$F$ $\max s_{G}^{ag}$ 0.02 0.04 0.04 0.05 0.06 0.14 0.33 0.01 0.06 0.14
PT $\max s_{G}^{c}$ 1.66 3.01 7.42 1.48 1.68 1.99 3.11 1.68 19.52 0.82
PT $\max s_{G}^{com}$ 0.11 0.18 0.24 0.08 0.15 0.01 0.66 0.08 0.16 0.14
P T $\max s_{G}^{ag}$ 0.00 0.10 0.14 0.06 0.11 0.03 0.28 0.10 0.02 0.56
Table 2: Cross‑lingual gender bias (max‑skew). For ten languages, we report the largest gender‑skew value ($s_{G}$) per stereotype axis—crime (c), communion (com), agency (ag)—on both FairFace($F$) and PATA (PT). Higher numbers indicate stronger association gaps; coloured entries mark the worst axis for each model, red for $\max s_{G}^{c}$, orange for $\max s_{G}^{\text{com}}$, and green for $\max s_{G}^{\text{ag}}$.
Model Data Metric en es fa fi fr hi pt sl tr xh
mCLIP $F$ $\mathrm{KL}_{\text{ag}}^{\mathrm{sym}}$ 0.00 0.01 0.03 0.03 0.03 0.14 0.01 0.02 0.00 0.01
$\mathrm{KL}_{\text{com}}^{\mathrm{sym}}$ 0.15 0.04 0.01 0.01 0.02 0.05 0.04 0.01 0.08 0.00
$\mathrm{KL}_{\text{c}}^{\mathrm{sym}}$ 0.02 0.02 0.01 0.00 0.02 0.00 0.01 0.06 0.01 0.00
PT $\mathrm{KL}_{\text{ag}}^{\mathrm{sym}}$ 0.00 0.00 0.06 0.01 0.00 0.01 0.01 0.04 0.00 0.11
$\mathrm{KL}_{\text{com}}^{\mathrm{sym}}$ 0.06 0.02 0.00 0.00 0.01 0.06 0.02 0.00 0.01 0.00
$\mathrm{KL}_{\text{c}}^{\mathrm{sym}}$ 0.02 0.02 0.01 0.00 0.02 0.00 0.01 0.06 0.01 0.00
NLLB $F$ $\mathrm{KL}_{\text{ag}}^{\mathrm{sym}}$ 0.02 0.03 0.08 0.19 0.02 0.03 0.02 0.09 0.06 0.05
$\mathrm{KL}_{\text{com}}^{\mathrm{sym}}$ 0.00 0.02 0.09 0.19 0.00 0.02 0.03 0.09 0.01 0.00
$\mathrm{KL}_{\text{c}}^{\mathrm{sym}}$ 0.03 0.03 0.00 0.03 0.02 0.03 0.02 0.04 0.08 0.06
PT $\mathrm{KL}_{\text{ag}}^{\mathrm{sym}}$ 0.07 0.02 0.04 0.18 0.03 0.05 0.05 0.04 0.07 0.15
$\mathrm{KL}_{\text{com}}^{\mathrm{sym}}$ 0.03 0.01 0.01 0.01 0.01 0.05 0.02 0.00 0.02 0.07
$\mathrm{KL}_{\text{c}}^{\mathrm{sym}}$ 0.00 0.00 0.02 0.03 0.02 0.03 0.02 0.04 0.06 0.05
CAPIVARA $F$ $\mathrm{KL}_{\text{ag}}^{\mathrm{sym}}$ 0.00 0.00 0.22 0.00 0.52 0.02 0.01 0.00 0.03 0.02
$\mathrm{KL}_{\text{com}}^{\mathrm{sym}}$ 0.05 0.02 0.01 0.00 0.01 0.03 0.01 0.04 0.09 0.40
$\mathrm{KL}_{\text{c}}^{\mathrm{sym}}$ 0.00 0.00 0.05 0.00 0.00 0.05 0.00 0.00 0.08 0.15
PT $\mathrm{KL}_{\text{ag}}^{\mathrm{sym}}$ 0.00 0.00 0.01 0.04 0.03 0.07 0.01 0.01 0.03 0.52
$\mathrm{KL}_{\text{com}}^{\mathrm{sym}}$ 0.00 0.00 0.01 0.01 0.01 0.05 0.00 0.04 0.02 0.05
$\mathrm{KL}_{\text{c}}^{\mathrm{sym}}$ 0.00 0.00 0.01 0.04 0.00 0.05 0.01 0.01 0.03 0.50
SIGLIP‑2 $F$ $\mathrm{KL}_{\text{ag}}^{\mathrm{sym}}$ 0.00 0.00 0.00 0.01 0.00 0.03 0.03 0.00 0.01 0.05
$\mathrm{KL}_{\text{com}}^{\mathrm{sym}}$ 0.08 0.03 0.00 0.00 0.00 0.03 0.04 0.07 0.01 0.00
$\mathrm{KL}_{\text{c}}^{\mathrm{sym}}$ 0.01 0.12 0.02 0.11 0.01 0.03 0.01 0.00 0.41 0.02
PT $\mathrm{KL}_{\text{ag}}^{\mathrm{sym}}$ 0.00 0.00 0.00 0.00 0.00 0.00 0.01 0.01 0.00 0.09
$\mathrm{KL}_{\text{com}}^{\mathrm{sym}}$ 0.00 0.00 0.03 0.00 0.01 0.00 0.06 0.01 0.01 0.01
$\mathrm{KL}_{\text{c}}^{\mathrm{sym}}$ 0.01 0.03 0.15 0.37 0.00 0.01 0.01 0.00 0.20 0.01
Table 3: Cross‑lingual gender bias (symmetric KL). Symmetric KL divergence ($\mathrm{SKL}$) between male‑ and female‑conditioned score distributions for each stereotype axis. Unlike max‑skew, SKL captures distributional shifts even when extreme outliers are absent.Higher numbers indicate more gender bias; coloured entries mark the worst axis for each model, red for $\mathrm{KL}_{\text{c}}^{\mathrm{sym}}$, orange for $\mathrm{KL}_{\text{com}}^{\mathrm{sym}}$, and green for $\mathrm{KL}_{\text{ag}}^{\mathrm{sym}}$.
Model Data Metric en es fa fi fr hi pt sl tr xh
mclip $F$ $\max s_{R}^{c}$ 0.40 0.40 0.27 0.10 0.39 0.08 0.26 0.41 0.23 0.04